Question 6 4 points Save Answer After an average acceleration of 3.09 m/s2 during 2.81 s, your car reaches a velocity of 145 m/s. Find the cars initial velocity O a-5.82 m/s O b. 0.00 m/s C 5.82 m/s d. 23.1 m/s
Answer #1

We know that a= vf-vi/t

a= acceleration= 3.09 m/s^2 vf=final velocity= 14.5 m/s vi= initial velocity= ? ; t= time = 2.81 s

so

3.09= 14.5-vi/2.81

vi=5.817 m/s = 5.82 m/s